A ball with mass of 0.2 kg collides with wall with a speed of 10 m/s at a right angle and rebounds with the same speed in the opposite direction. If the contact time between the ball and the wall is 0.1 s. What is the force exerted on the ball by the wall?


1
Expert's answer
2022-03-23T13:52:56-0400

The Newton's second law says

"F=\frac{m\Delta v}{\Delta t}=\frac{0.2\times (10-(-10))}{0.1}=40\:\rm N"